OpenStudy (anonymous):

simplify sqrt300n^11

OpenStudy (anonymous):

\[\sqrt{300n^{11}}\] ?

OpenStudy (anonymous):

yup. that's it.

OpenStudy (anonymous):

factorize it

OpenStudy (anonymous):

\(\large \sqrt{300n^{11}}=\sqrt{3\cdot 100 \cdot (n^5)^2 \cdot n} \) now simplify.

OpenStudy (anonymous):

how do you do the variable?